# DesignRepair: Dual-Stream Design Guideline-Aware Frontend Repair with Large Language Models

### v1

@@ IV-B1 Experimental Setup
UIS-Hunter [19] represents a significant effort in design violation detection by encompassing 23 types of UI components against 126 visual guidelines. In comparison, our tool collects 399 guidelines for 24 UI component types, including both do/don’t and general guidelines. Due to the substantial overlap in component-level hard constraint guidelines, and since UIS-Hunter primarily focuses on do/don’t guidelines (referred to as hard constraints), we selected UIS-Hunter as our baseline for detection performance comparison.
For a fair comparison to UIS-Hunter, we selected corresponding instances involving hard constraint violations: 127 issues from Vercel’s V0 and 48 issues from GitHub. We then measured the precision and recall of UIS-Hunter and our tool, comparing their outputs with manually assigned ground truth. Successful detection was defined as correctly identifying faulty elements. Table I shows the precision and recall comparison between our DesignRepair and UIS-Hunter.
After reviewing the UIS-Hunter project code, we identified limitations in its rule-based method. Specifically, it struggled with component mapping. Firstly, elements like a button within a navigation bar couldn’t be classified under multiple types. Additionally, UIS-Hunter was designed for Android apps and couldn’t handle the complexity of web components used in frameworks like React and Tailwind CSS.
To address these issues and enable UIS-Hunter to work on the evaluation UI data, we focused on comparing guideline violation detection performance. Two of our authors manually mapped component names to their corresponding types, which took about one hour per project. While our method uses LLMs for automated component analysis and mapping, demonstrating greater flexibility than the rule-based approach.
@@ IV-B2 Results
The recall rates, as shown in Table I, indicate that UIS-Hunter achieves 32.3% for Vercel’s V0 projects and 22.9% for GitHub projects, while our method achieves 92.1% and 83.3%, respectively.
UIS-Hunter’s lower true positive rates are caused by its fixed rules and thresholds, which miss many issues. It cannot assess active state components, handle different window sizes, or detect small component issues like missing button semantics. Specifically, it cannot assess the correctness of components in active states and only identifies issues in static screenshots through image processing. Moreover, UIS-Hunter fails to detect issues that arise from incorrect code in different window sizes, such as line breaks not displaying in larger windows, which would require multiple detections (need high time costs in testing). Additionally, it cannot identify missing button semantics or detect small components. For instance, UIS-Hunter cannot distinguish and evaluate the semantic differences between buttons for “follow” and “message”, which require different active colors based on their semantics.
In contrast, our algorithm overcomes these limitations, ensuring high recall and consistency detection among components, such as uniform font colors across multiple buttons. It also accurately identifies minor visual violations, such as padding differences between the left and right, which visual methods might miss.
For precision, as shown in Table I, UIS-Hunter achieves 55.4% for Vercel’s V0 projects and 47.8% for GitHub projects, while our method achieves 87.3% and 90.9%, respectively. UIS-Hunter’s lower precision is largely due to false positives caused by OCR errors, such as spaces in long text strings. Another cause is the fixed nature of rule-based approaches, such as requiring colored dividers in lists or limiting the number of icons, which also contributes to the higher false positive rate.
@@
IV-C RQ2: What is the contribution of each key repair strategy to the overall performance of our approach?
@@ IV-C1 Experimental Setup
To evaluate our DesignRepair’s effectiveness in correcting guideline violations for frontend code, we conducted a step-by-step analysis. This included Component Extraction, Component Mapping, Violation Detection, and different aspects of Violation Repair, including overall repair.
In the evaluation, we compared outputs with manually assigned ground truth for each step. ”Component Extraction” was successful if the component’s name was found in the file, while ”Component Mapping” was successful if the component type was correctly identified in the material design component list. Successful detection and repair were defined as correctly identifying and fixing faulty elements based on the guidelines. Table II shows the precision and recall comparison.
To separately evaluate the effectiveness of component/property and soft/hard constraint strategies, we tested each strategy in isolation by checking related guideline subsets and evaluating accuracy and precision as shown in IV-A3. For example, for component-only violation repair, we disabled the property repair module and calculated accuracy and recall only for component-specific guidelines.
@@ IV-C2 Results
Component Extraction Results: Upon examining DesignRepair’s performance in Component Extraction, we identified 88 component types across 20 Vercel’s V0 projects and 64 component types for GitHub projects. As shown in II, Vercel’s V0 projects achieved 96.7% recall and 93.6% precision, while GitHub projects achieved 96.5% recall and 91.8% precision. The slightly lower precision compared to recall is due to the LLM extracting more components when uncertain, to avoid missing any. We also found that the missed components had minimal impact on mapping performance since they were often the same type as other correctly identified components, differing only in specific names.
@@
Component Mapping Results: For Component Mapping, we identified 88 component types in 51 Vercel’s V0 projects and 39 component types in GitHub projects. As shown in II, Vercel’s V0 projects achieved 94.1% recall and 90.5% precision, while GitHub projects achieved 94.9% recall and 90.2% precision.
@@
Complex components formed from simpler ones, such as navigation bars, cards, and lists, achieved nearly 100% accuracy and recall. The LLM effectively analyzed and identified these groups, which is challenging for rule-based methods. However, simple components defined primarily by CSS, like badges and dividers, were not well-extracted by our method.
@@
In general, our DesignRepair results demonstrate that the LLM-based analysis method can handle complex component extraction tasks, even when frontend code often prioritizes appearance over functionality, with some button components represented by hyperlinks instead of proper button tags. Rule-based approaches struggle to correctly identify such variations.
@@
Violation Detection Results: We evaluated guideline violations for 196 Vercel’s V0 and 115 GitHub’s. As shown in II, Vercel’s V0 projects achieved 92.9% recall and 90.1% precision, while GitHub projects achieved 87.8% recall and 93.5% precision. Our method effectively detects both component and property issues. For components, it identifies violations in anatomy, behavior, layout, usage, and placement. For properties, it detects issues related to labels, text, color, and spacing, focusing on accessibility problems like alt-text and labeling.
When comparing Vercel’s V0 and GitHub projects, LLM-generated projects recall more candidates for repair. However, when more complex pages are detected and repaired simultaneously, recall decreases due to LLMs avoiding large adjustments to prevent new issues. Furthermore, longer related guidelines and code text lengths also lower the LLM recall rates for GitHub projects.
@@
Component/Property Only Repair Results: For component-only repair, we evaluated 149 Vercel’s V0 violations instances and 60 GitHub’s. As shown in II, Vercel’s V0 achieved 89.9% recall and 87% precision, while GitHub achieved 81.7% recall and 87.5% precision. For property-only repair, we evaluated 47 Vercel’s V0 violations instances and 55 GitHub’s. Vercel’s V0 achieved 87.2% recall and 85.4% precision, while GitHub achieved 89% recall and 94.2% precision.
@@
Soft/Hard Constraint Only Repair Results: For soft constraint-only repair, we evaluated 27 Vercel’s V0 violations instances and 15 GitHub’s. Vercel’s V0 achieved 70.4% recall and 82.6% precision, while GitHub achieved 67.7% recall and 77% precision. For hard constraint-only repair, we evaluated 169 Vercel’s V0 projects and 100 GitHub projects. Vercel’s V0 achieved 92.3% recall and 87.2% precision, while GitHub achieved 88% recall and 90.7% precision.
The performance of soft constraint repair is generally lower than that of hard constraint repair and Hard constraint repairs on GitHub had lower recall but improved precision. Those are due to the conservative approach of LLMs to avoid introducing new issues in complex pages, resulting in fewer true positives and false negatives, if. Vercel’s V0 projects showed better performance on soft constraints, highlighting the LLM’s ability to enhance design diversity and aesthetics when hard constraint influence decreases.
Overall Repair Results: For overall repair, Vercel’s V0 projects achieved 89.3% recall and 86.6% precision, while GitHub projects achieved 85.2% recall and 90.7% precision.
Our approach effectively adheres to design guidelines without extensive training. For instance, as illustrated in Fig. 6(a) and Fig. 6(b), DesignRepair adeptly solves element overlapping issue and enhances background color contrast, making the design accessibility-friendly and visually appealing. Furthermore, as showcased in Fig. 6(c) and Fig. 6(d), our tool effectively addresses responsive layout issues, ensuring optimal appearance across various window sizes. Another striking example, presented in Fig. 1(b), demonstrates our software’s proficiency in unifying component colors and improving aspects like accessibility, text legibility, and contrast.
@@
IV-D RQ3: What is the perceived quality of the fixes generated by our approach from a user perspective?
@@ IV-D1 Procedure
We recruited a total of 26 participants (12 female,14 male), aged between 25 and 35. The group comprised 6 researchers, 10 students, and 10 software developers from diverse countries including Australia 8, China 15, and the United States 3. The educational backgrounds of the participants were varied: 5 had Bachelor’s degrees, 15 had Master’s degrees, and 6 had Ph.D.s. Their affiliations were diverse, with 6 from a science agency, 10 from major internet companies, and 10 from universities.
@@ IV-D2 Results
Knowing the order of the videos, we calculated how often the before UI page scored lower, equal to, or higher than the after-repair UI page, and drew Fig. 8(b), which offer insightful conclusions regarding user perceptions of the repaired pages. A notable 58.7% of users regarded the pages as visually superior before-repair. In the interview, participants provided feedback indicating an enhanced user experience, evidenced by greater ease of use, more efficient information retrieval, and a noticeable increase in design attractiveness.
Interestingly, 20.4% of participants rated the before and after repair versions equally. Feedback from these participants’s interview indicated that in cases where the frontend code was complex and the errors minor (involving aspects like content and accessibility features), the distinctions in repair quality were not immediately apparent. Conversely, 20.1% of participants expressed a preference for the original design. After being informed about the guideline checks and repairs for readability and interactivity, some participants still preferred the original versions. They felt that certain soft constraint repairs were subjective. Moreover, some design issues, like colorful themes, did not affect usability and were more engaging.
In conclusion, our approach demonstrates its effectiveness through both statistical and descriptive analyses.
@@ V Discussion and Future Work
Our work presents an extensible, knowledge-based code repair method that effectively utilizes design guidelines without requiring manual rule building and development. This approach’s strength lies in its flexibility to support and integrate diverse guidelines. By leveraging LLMs and an extendable knowledge-driven framework, our method can adapt seamlessly to evolving guidelines. Furthermore, future improvements could focus on mechanisms to effortlessly incorporate guideline updates and further narrow the scope of LLM inspection. We have identified a promising direction by combining related specifications with relevant code, properties, and values to achieve more effective performance. We found that shortening the reasoning process can enhance code repair capabilities. It is more efficient to have LLMs and traditional analysis tools, such as render page analyzing tool, cooperate and leverage their respective advantages. These advancements would ensure continued alignment with the latest industry and company standards, allowing organizations to easily tailor the approach to their specific needs.
Additionally, our work provides a first inspection of LLM-generated frontend code issues and methods to improve them, which is valuable for enhancing generated code, especially frontend code. In the future, we can incorporate LLM issue detection and repair during the code generation process, enabling early problem identification and the creation of more reliable code.
@@ VI-A Accessibility Testing Tools
Widely-used accessibility testing tools like Playwright [23], Google Lighthouse [24], and axe-core [25] are primarily designed for addressing accessibility issues like text contrast, icon tapability, image accessibility, display issues, and intention-practice mismatches. However, these tools lack comprehensive repair capabilities and often fail to detect a significant portion of issues. A user study [36] revealed that these tools only reported around 50% of the issues they claimed to detect, highlighting their limitations in providing comprehensive QA solutions.
@@ VI-B Research for UI Issues
To address the limitations of accessibility testing tools, researchers have developed tools targeting specific UI issues. OwlEye [16], IFIX [37], CBRepair [38], and MFIX [37] focus on Android GUI layout issues, web application internationalization presentation problems, and mobile web presentation issues, respectively. Iris [39] and Seenomaly [18] are specifically designed for color-related accessibility issues and detecting violations of design guidelines in animations. While these tools have contributed to advancing QA in their respective domains, they lack a more general, adaptive, and effective ways to address the diverse range of design issues in frontend code.
UIS-Hunter [19] represents a significant effort in design smell detection by encompassing 23 types of UI components against 126 visual guidelines. This extensive scope marks a notable advancement in identifying a wide array of design issues. However, despite its broad coverage, UIS-Hunter’s approach has limitations. It offers a broader spectrum of issues to detect compared to tools focused on specific issues, but it still falls short in analyzing complex high-level design principles. UIS-Hunter primarily detects violations of guidelines for individual components but fails to offer insights into higher-level aspects, such as the intricate layout effects between components and theme color consistency across the application. Furthermore, it can only detect issues and lacks repair ability.
@@ VI-C LLM-based Code Generation and Repair
The emergence of automated program repair research, coupled with the advancements in LLMs, opens up new possibilities for an open approach for program repair. Pre-trained LLMs like CodenBERT [40], CodeX [41], AlphaCode [42], and Code Llama [2] have demonstrated significant improvements in code generation and repair. Research on performance enhancement in program repair has also progressed, with models like PLBART [43], CodeT5 [44], CodeGen [45], and InCoder [46] showing competitive or superior performance compared to deep learning-based techniques. These models leverage knowledge-driven methods to enhance repair capabilities without additional model training.

### v2

@@ IV-B1 Experimental Setup
We used UIS-Hunter [19] as our baseline, as it also focuses on detecting design violations against the material design. It primarily focuses on do/don’t guidelines (i.e., hard constraints), and only supports 23 types of UI components and examines 126 visual design guidelines. In comparison, our tool collects 399 guidelines for 24 UI component types, including both do/don’t and general guidelines.
To make a fair comparison, we selected corresponding instances involving hard constraint violations: 127 issues from Vercel’s V0 and 48 issues from GitHub, to perform the experiment. We adopted the precision and recall metrics. A detection result is considered as a true positive if the elements are actually faulty.
During the experiment, we identified limitations in UIS-Hunter due to its rule-based approach, which struggles with component mapping. For instance, elements like buttons within a navigation bar should be classified under multiple types, but UIS-Hunter restricts them to a single type. Moreover, UIS-Hunter was specifically designed for Android apps and lacks the capability to handle the complexity of web components commonly used in frameworks such as React and Tailwind CSS. To address these issues and adapt UIS-Hunter for the experimental data (as some component names in the experimental data differ from those supported in UIS-Hunter), we focused on evaluating its performance in detecting guideline violations. Two authors manually mapped component names to their corresponding types, which took about one hour per project. In contrast, our method enables automated component analysis and mapping, demonstrating significantly greater flexibility than the rule-based approach.
@@ IV-B2 Results
As shown in Table I, for recall, UIS-Hunter achieves 32.3% for Vercel’s V0 projects and 22.9% for GitHub projects, while our method achieves 92.1% and 83.3%, respectively.
UIS-Hunter’s lower true positive rates stem from its reliance on fixed rules and thresholds, which result in many issues being overlooked. We identified three primary failure reasons for UIS-Hunter. First, it is unable to evaluate active state components. It cannot assess the correctness of components in active states and can only identify issues in static screenshots through image processing. Second, UIS-Hunter fails to detect responsive design issues such as line breaks not displaying correctly in larger windows. Addressing these issues would require multiple detections, a capability that this method lacks. Moreover, it cannot capture the semantics of components, resulting in an inability to detect subtle component issues. For instance, UIS-Hunter cannot distinguish and evaluate the semantic differences between buttons for “follow” and “message”, which require different active colors based on their semantics.
In contrast, our algorithm overcomes these limitations, ensuring high recalls and consistency detection among components, such as ensuring uniform font colors across multiple buttons. It also accurately identifies minor visual violations, such as padding differences between the left and right, which purely visual-based methods may miss.
For precision, as shown in Table I, UIS-Hunter achieves 55.4% for Vercel’s V0 projects and 47.8% for GitHub projects, while our method achieves 87.3% and 90.9%, respectively. UIS-Hunter’s lower precision is largely due to false positives caused by OCR errors, such as spaces in long text strings. Additionally, its rigid rule-based approach-such as requiring colored dividers in lists or limiting the number of icons-further contributes to the higher false positive rate.
@@
IV-C RQ2: Effectiveness in repairing violations
@@ IV-C1 Experimental Setup
To evaluate the effectiveness of DesignRepair in correcting guideline violations in frontend code, we conducted an assessment of each step module. This evaluation includes analyzing Component Extraction, Component Mapping, Violation Detection, various aspects of Violation Repair, and the overall performance of the tool.
In the evaluation, we compared outputs with manually annotated ground truth for each step. For ”Component Extraction”, an extraction is deemed correct if the component name can be found in the file. For ”Component Mapping”, a mapping is considered as correct if the mapped type matches an entry in the Material Design component list. Successful detection and repair are defined as correctly identifying and fixing faulty elements based on the guidelines.
To evaluate the effectiveness of component/property and soft/hard constraint strategies separately, we tested each strategy in isolation and evaluate them on their respective guideline subsets. For example, for component-only violation repair strategy, we disabled the property repair module and calculated results only for component-specific guidelines.
@@ IV-C2 Results
Component Extraction Results: Upon examining DesignRepair’s performance in Component Extraction, we identified 88 component types across 20 Vercel’s V0 projects and 64 component types for GitHub projects. As shown in II, DesignRepair achieved 96.7% recall and 93.6% precision on Vercel’s V0 projects , and 96.5% recall and 91.8% precision on GitHub projects. Precisions are slightly lower than recalls because the LLM used an aggressive strategy to extract more components when uncertain, minimizing omissions. We found that the missed components in our detection had little impact on mapping performance because they were usually the same type as the correctly identified components, such as “buttons”, differing only in their specific names, like “likebutton” or “sharebutton”.
@@
Component Mapping Results: We identified 88 component types in 51 Vercel’s V0 projects and 39 component types in GitHub projects. As shown in II, our model achieved 94.1% recall and 90.5% precision on Vercel’s V0 projects, and 94.9% recall and 90.2% precision on GitHub projects.
@@
Complex components formed from simpler ones, such as navigation bars, cards and lists, achieved nearly 100% precision and recall. The LLM effectively analyzed and identified these groups, which is challenging for rule-based methods. However, simple components defined primarily by CSS, like badges and dividers, were less effectively extracted by our method.
@@
In general, the results demonstrate that the LLM-based analysis method can handle complex component extraction tasks and can deal with various, sometimes anecdotal implementation behaviors, such as using hyperlinks instead of proper button tags to represent a button. Rule-based approaches struggle to correctly identify such variations.
@@
Violation Detection Results: We evaluated guideline violations on 196 Vercel’s V0 and 115 GitHub’s. As shown in Table II, DesignRepair achieved 92.9% recall and 90.1% precision on Vercel’s V0 projects, and 87.8% recall and 93.5% precision on GitHub projects. Our method effectively detects both component and property issues. For components, it identifies violations in anatomy, behavior, layout, usage, and placement. For properties, it detects issues related to labels, text, color, and spacing, focusing on accessibility problems like alt-text and labeling.
In our analysis, our method demonstrates higher recall of candidate fixes on Vercel’s V0 compared to GitHub projects. However, recall decreases on complex GitHub pages as DesignRepair’s LLM avoids significant adjustments to prevent introducing new issues.
@@
Component/Property Only Repair Results: For component-only repair, we evaluated 149 Vercel’s V0 violations instances and 60 GitHub’s. As shown in Table II, our method achieved 89.9% recall and 87% precision on Vercel’s V0, and 81.7% recall and 87.5% precision on GitHub projects. For property-only repair, we evaluated 47 Vercel’s V0 violations instances and 55 GitHub’s. Our model achieved 87.2% recall and 85.4% precision on Vercel’s V0, and 89% recall and 94.2% precision on GitHub projects.
@@
Soft/Hard Constraint Only Repair Results: For soft constraint-only repair, we evaluated 27 Vercel’s V0 violations instances and 15 GitHub’s. Our model achieved 70.4% recall and 82.6% precision on Vercel’s V0 projects, and 67.7% recall and 77% precision on GitHub projects. For hard constraint-only repair, we evaluated 169 Vercel’s V0 projects and 100 GitHub projects. The results show 92.3% recall and 87.2% precision for Vercel’s V0, and 88% recall and 90.7% precision for GitHub projects.
The performance of soft constraint repair is generally lower than that of hard constraint repair. On GitHub project, results for hard constraint repairs have lower recall but improved precision. These are due to the conservative approach of LLMs to avoid introducing new issues in complex pages, resulting in fewer true positives and false negatives. On Vercel’s V0 projects, soft constraint repairs perform better, showcasing the LLM’s ability to enhance design diversity and aesthetics when the tested pages have fewer hard constraint issues.
@@
Overall Repair Results: For overall repair, the proposed method achieved 89.3% recall and 86.6% precision on Vercel’s V0 projects, and 85.2% recall and 90.7% precision on GitHub projects. Our approach effectively adheres to design guidelines without extensive training. For instance, as illustrated in Fig. 6(a) and Fig. 6(b), DesignRepair adeptly solves element overlapping issue and enhances background color contrast, making the design accessibility-friendly and visually appealing. Furthermore, as showcased in Fig. 6(c) and Fig. 6(d), our tool effectively addresses responsive layout issues, ensuring optimal appearance across various window sizes. Another striking example, presented in Fig. 1(b), demonstrates our software’s proficiency in unifying component colors and improving aspects like accessibility, text legibility, and contrast.
IV-D RQ3: Perceived usefulness
@@ IV-D1 Procedure
We recruited a total of 26 participants (12 female,14 male), aged between 25 and 35. The group comprised 6 researchers, 10 students, and 10 software developers from diverse countries including Australia 8, China 15, and the United States 3. The educational backgrounds of the participants were varied: 5 had Bachelor’s degrees, 15 had Master’s degrees, and 6 had Ph.D.s. Their affiliations were diverse, with 6 from science agencies, 10 from major internet companies, and 10 from universities.
@@ IV-D2 Results
Knowing the order of the videos, we calculated how often the before UI page scored lower, equal to, or higher than the after-repair UI page. As shown in Fig. 8(b), a notable 58.7% of users regarded the pages as visually superior before-repair. In the interview, participants provided feedback indicating an enhanced user experience, evidenced by greater ease of use, more efficient information retrieval, and a noticeable increase in design attractiveness.
Interestingly, 20.4% of participants rated the before and after repair versions equally. Feedback from these participants’s interview indicated that in cases where the frontend code was complex and the errors minor (involving aspects like content and accessibility features), the distinctions in repair quality were not immediately apparent. Conversely, 20.1% of participants expressed a preference for the original design. After being informed about the guideline checks and repairs for readability and interactivity, some participants still preferred the original versions. They felt that certain soft constraint repairs were subjective. Moreover, some design issues, like colorful themes, did not affect usability and were more engaging. Overall, our approach demonstrates its effectiveness through both statistical and descriptive analyses.
@@ V Discussion and Future Work
Our work presents an extensible, knowledge-based code repair method that effectively utilizes design guidelines without requiring manual rule building and development. This approach’s strength lies in its flexibility to support and integrate diverse guidelines. By leveraging LLMs and an extendable knowledge-driven framework, our method can adapt seamlessly to evolving guidelines. Furthermore, future improvements could focus on mechanisms to effortlessly incorporate guideline updates and further narrow the scope of LLM inspection. We have identified a promising direction by combining related specifications with relevant code, properties, and values to achieve more effective performance. We found that shortening the reasoning process can enhance code repair capabilities. Combining LLMs with traditional page analysis tools can create a more efficient workflow by leveraging the unique strengths of each tools. These advancements ensure alignment with industry standards, enabling organizations to tailor the approach to their specific needs.
Additionally, our work provides a first inspection of LLM-generated frontend code issues and methods to improve them, which is valuable for enhancing generated code, especially frontend code. In the future, we can incorporate LLM issue detection and repair during the code generation process, enabling early problem identification and more reliable code creation.
@@ VI-A Accessibility Testing Tools
Widely-used accessibility testing tools like Playwright [23], Google Lighthouse [24], and axe-core [25] are primarily designed for addressing accessibility issues like text contrast, icon tapability, image accessibility, display issues, and intention-practice mismatches. These tools lack various aspects of repair capabilities and often miss many issues. A user study [35] found they detected only about 50% of the issues they claimed, highlighting their limitations.
@@ VI-B Research for UI Issues
To address the limitations of accessibility testing tools, researchers have developed tools targeting specific UI issues. OwlEye [16], IFIX [36], CBRepair [37], and MFIX [36] focus on Android GUI layout issues, web application internationalization presentation problems, and mobile web presentation issues, respectively. Iris [38] and Seenomaly [18] are specifically designed for color-related accessibility issues and detecting violations of design guidelines in animations. While these tools have contributed to advancing QA in their respective domains, they lack general, adaptive, and effective methods to address diverse frontend design issues.
UIS-Hunter [19] represents a vital effort in design smell detection by encompassing 23 types of UI components against 126 visual guidelines. This extensive scope marks a notable advancement in identifying a wide array of design issues. Despite its broad coverage, UIS-Hunter struggles with complex design principles, focusing on component-level guidelines while missing higher-level aspects like layout issues. Furthermore, it can only detect issues and lacks repair ability.
@@ VI-C LLM-based Code Generation and Repair
The emergence of automated program repair research, coupled with the advancements in LLMs, opens up new possibilities for an open approach for program repair. Pre-trained LLMs like CodenBERT [39], CodeX [40], AlphaCode [41], and Code Llama [2] have demonstrated significant improvements in code generation and repair. Research on performance enhancement in program repair has also progressed, with models like CodeT5 [42], CodeGen [43], and InCoder [44] showing competitive or superior performance compared to deep learning-based techniques. These models leverage knowledge-driven methods to enhance repair capabilities without additional model training.